OpenStudy (anonymous):

Check my answer? What is the solution of 3|x+5|≤6? x ≤ -3 and x ≥ -7 No solution x ≥ -3 and x ≤ -7 x < -3 or x > -7

OpenStudy (anonymous):

My answer is A

OpenStudy (anonymous):

The |'s are ( and ) right?

OpenStudy (anonymous):

I'm not sure actually, that is what threw me off.

OpenStudy (anonymous):

Correct!

OpenStudy (anonymous):

The question actually shows |?

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

Yeah it did

OpenStudy (anonymous):

3 | x + 5 | <= 6 --- divide by 3 |x + 5 | <= 2 -2 <= x + 5 <= 2 -- subtract 5 -2 - 5 <= x <= 2 - 5 -7 <= x <= -3 yep...your right

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

sure thing :)